How much water must be added to 6gallons of a 4% insecticide solution to reduce the concentration to 3% (hint: water is 0% insecticide)
Answer by HyperBrain(694)

Let x=amount of water required
---
6 gallons + x gallons = (x+6) gallons of mixture
For x gallons of water, we have 0 gallon of insecticide
For 6 gallons of 4% insecticide solution, we have 0.04(6)=0.24 gallon of insecticide
---
We have the equation
0.03(x+6)=0.24
3(x+6)=24
x+6=8
x=2 gallons